• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Zhihong Ye

    Log entry time 03:11:30 on March 04, 2012

    Entry number 363304

    keyword=G2p Status Update

    I am summarizing major activities performed during Saturday 8am and
    Sunday Owl Shift:
    1, BPM & Harp Scan
       Downstream BPM X- & Y- have double peaks. John Musson came in at 9am
    and explained that he saw large noisy background exist in those channels
    can cause the issue if we run at low current and turn on auto-gain. He
    will diagonalize the issue on Monday.
       Harp Analyzer was reported to be fixed but when Alex and Pengjia tried
    to do Harp scan at 9pm, they could not see Harp signals. So far we have
    not found out what exactly the issue is. We will call for experts on
    Sunday day time for help.
       We currently turned on raster, asked for 50nA and checked Spot++ with
    cross hair but it barely could be seen. JP, Alex, Toby and Pengjia worked
    together to manually move the C12 target to a best location where beam
    should hit roughly near the center of the target. We should be able to
    safely take optics data with C12 target.
    2, Moller Measurement
       Sasha and his folks started Moller Measurement on 2pm and finished it
    on 8pm Saturday. Detail results can be checked in :
    3, Large Asym Measreument
       Kalyan and Chao worked on large asym measurement from 6pm together
    with Moller measurement. 
    4, Two BCM Tungsten Measurement
       Alex did two Tungsten measurement at 12pm and 8pm Saturday.
    5, RHRS-S1-TDC double peaks
       Ryan and Melissa found out the strange 500ns wide double peaks in the
    first 7 TDC channels. They went  down the hall to check the connection
    and issue looked gone.
    6, RHRS-VDC-1 Low Efficiencies
       RHRS VDCs tripped since Friday 11pm and were fistly noticed at 7am. It
    was realized that the rate was very high at 5uA when they send the tune
    beam in the hall. We should keep eyes on the data rates and make sure HVs
    of VDCs are off if we have to run at extreme high rate.
       Low efficiency of VDC-U1&V1 planes was found at Saturday 5pm after we
    turned on HVs again and checked with beam on. The Max Hit Number of V1&U1
    is equal to 3 instead of 5, which indicates hardware issues. We are still
    checking what are the exact problems that cause the low efficiencies.
    7, Optics Data:
       Jixie, Min and Chao are preparing the optics data taking but we need
    to resolve R-VDC1 low efficiency issues.
    7, Control Access
       8am for Pengjia to disconnect cables that sending BPM signals to Happex.
       8pm for Ryan, Melissa to check R-S1-TDC issue, and Kalyan and Chao to
    work on Third Arm DAQ.
    9, Compton Work:
       It looks like we are not able to do the Compton Measurement before
    Monday down.

    A copy of this log entry has been emailed to: jpchen,slifer,rom,vasulk,camsonne,g2p